the period of a simple pendulum of length 80cm was found to have doubled when the length of the pendulum was increades by x. calculate x


1
Expert's answer
2021-06-21T11:38:22-0400

"T=2\pi \sqrt{\frac{l}{g}},"

"2T=2\pi \sqrt{\frac{l+x}{g}},"

"2\sqrt{\frac{l}{g}}= \sqrt{\frac{l+x}{g}},"

"4l=l+x,"

"x=3l,"

"x=3\cdot 80=240~cm."